OpenStudy (anonymous):

how do I solve this equation: logcuberoot(x^2+21x) = 2/3?

OpenStudy (anonymous):

\[\log \sqrt[3]{x ^{2}+21x} = 2/3\]

OpenStudy (lgbasallote):

a good idea is to change to exponential form first...

Parth (parthkohli):

Do you know the definition of log?\[\log_a b = c \implies a^c =b\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Logarithms are exponents, so the 2/3 is the exponent on the base of the logarithm to get the power cuberoot(x^2+21x)

OpenStudy (anonymous):

Hint: This can be rewritten as 3x^2+63x-2=0. Can this be solved using the quadratic equation?

OpenStudy (anonymous):

oops i did not see the log.

OpenStudy (anonymous):

Hint: power rule then 10^2